1. If the inverse demand curve of profit maximizing monopolist is given as P =1200 − 2Q , and cost function as

 C = Q3 − 61.25Q2+1528.5Q + 2000, find equilibrium output level, monopolist price, and profit. (6Marks)



1
Expert's answer
2021-07-05T17:36:45-0400

"P = 1200-2Q"

"C= Q^3-61.25Q^2+ 1528.5Q + 2000"

Equilibrium output occurs when "MR= MC"

Total Revenue"= P\u00d7 Q"

= "(1200-2Q)Q = 1200Q - 2Q^2"

Marginal Revenue MR is the first derivative of Total Revenue TR

"\frac{dTR}{DQ}" "= 1200- 4Q"

Marginal Cost, MC is

"\frac{DC}{DQ} = 3Q^2- 122.5Q+1528.5"

At equilibrium

MR= MC

1200- 4Q = "3Q^2 - 122.5Q +1528.5"

"Q = 36.5"

Monopolist Price is;

"P =1200-2(36.5)= 1127"


Profit = TR- TC

TR="1200(36.5)-2(36.5^2)= 40535.5"


TC = "24817.06" ( using Cost function with Q= 36.5)


Profit ="40535- 24817.06= 15717.94"